# Sequential Diversification with Provable Guarantees

### v2

@@ Abstract.
journalyear: 2025copyright: rightsretainedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanybooktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanyisbn: 979-8-4007-1329-3/25/03ccs: Information systems Information retrievalccs: Theory of computation Approximation algorithms analysis
@@ 1. Introduction
The model needs to be instantiated with a diversity measure for item sets, and any standard measure can be used. In this paper, we consider two commonly-used diversity measures: pair-wise sum diversity (Borodin et al., 2012) and coverage diversity (Ashkan et al., 2015; Puthiya et al., 2016; Zhai et al., 2015). For coverage diversity, we prove the problem is ordered-submodular (Kleinberg et al., 2023), and thus, existing techniques can be applied.
@@
Depending on the values of the continuation probabilities, we introduce two distinct algorithms. When continuation probabilities are relatively small, e.g., constant values that are smaller than $1$, we design an algorithm that focuses on selecting the top-$\tau$ items, where the value of $\tau$ controls the trade-off between runtime and approximation quality. Notably, when $\tau=2$, our approach aligns with a simple greedy approach, and hence, we provide the approximation ratio for this greedy approach. In contrast, when continuation probabilities are arbitrarily close to $1$, it’s not just the first few items that affect the sequential diversity score; rather, more items ranked toward the front impact the score. To address this case, we propose a second algorithm using a greedy matching approach on the set of all items, and offers a constant approximation guarantee.
In our experiments, we introduce metrics for user satisfaction and engagement. We experimentally compare our methods against baselines and demonstrate their effectiveness in finding high-quality solutions. Due to space limitations, proofs of our theoretical results and a more extensive experimental evaluation are provided in the appendix.
@@ 3.1. Problem definition
We will work with two commonly-used diversity functions. The first is the pair-wise sum diversity: for a distance function $d\colon U\times U\rightarrow\mathbb{R}_{\geq 0}$, the diversity function $\mathcal{D}_{+}(A)=\sum_{i,j\in A}d(i,j)$ sums all pair-wise distances of items in $U$.
@@
(3.1) $\mathcal{S}_{+}(O)=\mathbb{E}_{A\sqsubseteq O}[\mathcal{D}_{+}(A)]=\mathbb{E}_{A\sqsubseteq O}\left[\sum\nolimits_{i,j\in A}d(i,j)\right].$
@@ 3.2. Reformulation and complexity
To simplify our calculations, we use the notation $p_{O_{i}}=\prod_{t=1}^{i}p_{\pi(t)}$ and ${}~{}d(\pi(i+1),O_{i})=\sum\nolimits_{t=1}^{i}d(\pi(i+1),\pi(t))$.
@@
$\mathcal{S}_{+}(O)=\sum\nolimits_{i=1}^{n-1}p_{O_{i+1}}d(\pi(i+1),O_{i}).$
@@
Let $U=\{u_{1},u_{2},u_{3}\}$, where $d(u_{1},u_{2})=0.3$, $d(u_{1},u_{3})=1$ and $d(u_{2},u_{3})=1$. Let $p_{u_{1}}=p_{u_{2}}=1$ and $p_{u_{3}}=0$. The sequential sum diversity scores differ according to the order of the items, specifically:
@@ 4. Sequential coverage diversity
where ||||| | denotes the concatenation of sequences and elements.
@@ 5. Ordered Hamiltonian path
We are given a finite set $U=\{1,\ldots,n\}$ of $n$ distinct items, a distance function $d(\cdot,\cdot)$, and probabilities $\{p_{1},\ldots,p_{n}\}$ assigned to each item $i\in U$. Let $O=(\pi(i))_{i=1}^{n}$ be the ordered sequence of items of $U$ according to an order $\pi$. The ordered Hamiltonian path, denoted by $\mathcal{H}$, is defined as
(5.1) $\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1)),$
@@
Note that the ordered Hamiltonian-path differs significantly from the classic Hamiltonian-path (Gurevich and Shelah, 1987), where the edge weights do not depend on the order in which edges are placed. In contrast, in the ordered Hamiltonian-path, the edge weights $d(\pi(i),\pi(i+1))$ are multiplied by a coefficient $W_{O_{i}}$, which depends on the order $\pi$ as well as on the probabilities of all the items in $U$.
@@
We are given a finite set $U=\{1,\ldots,n\}$ of $n$ distinct items, a distance function $d(\cdot,\cdot)$, and probabilities $p_{1},\ldots,p_{n}$ assigned to each item $i\in U$. The goal is to find an order $O^{*}$ of the items in $U$ so as to maximize the $\mathcal{H}$ objective, that is,
@@
Let $p_{i}\in[a,b]$, with $10<a<b<$ 1. An $\alpha$-approximation solution for MaxOHP is a $\frac{a(1-b)}{2b(1-a)}\alpha$-approximation for MaxSSD.
Corollary 0.
Consider the special case where $p_{i}=p$ for all $i\in U$. An $\alpha$-approximation solution for MaxOHP is a $\frac{\alpha}{2}$-approximation for MaxSSD.
@@ 6. Uniform continuation probabilities
We analyze three regimes for $p$, and design algorithms that achieve constant-factor approximations for MaxSSD in each regime. Due to space constraints, two regimes are presented in the main content, with the third deferred to the appendix.
As outlined earlier, we reformulate the problem using the intemediate problem MaxOHP, which then leads to a solution for MaxSSD. Recall that for an order $O$, the ordered Hamiltonian-path objective is defined as $\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d(\pi(i),\pi(i+1))$, where $W_{O_{i}}=\sum_{j=i+1}^{n}p_{O_{j}}$. In the uniform probability setting, since $p_{i}=p$ for all $i$, the coefficient $W_{O_{i}}$ is entirely determined by its position in the order, rather than the actual items being ranked. In this case,
@@ 6.1. The best-$\tau$ items algorithm
To obtain a better intuition for this regime, consider a scenario where $p$ is very small, i.e., close to 00. In this case, a user is likely to quit after examining only a few items, so the ranking of the first few items becomes critical. To maximize diversity, the most diverse items should be placed at the top, as they contribute most significantly to the objective of the ordered Hamiltonian path.
Building on this insight, we propose an algorithm for the general case where $p$ is bounded away from 1. We call this algorithm best-$\tau$ items ($\textsc{B}\tau\textsc{I}$), where $\tau$ is a parameter controlling the approximation quality. The algorithm always outputs a ranking of all items, with the pseudocode provided in Algorithm 1.
@@
Algorithm 1 Best-$\tau$ items ($\textsc{B}\tau\textsc{I}$) algorithm
Concretely, the best-$\tau$ items ($\textsc{B}\tau\textsc{I}$) algorithm works as follows: for any integer $\tau$, let
^$\hat{\mathcal{H}}(O_{\tau})=\sum\nolimits_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d(\pi(i),\pi(i+1))$
denote the contribution of $O_{\tau}$ to $\mathcal{H}(O)$. The $\textsc{B}\tau\textsc{I}$ algorithm seeks the optimal $\tau$-item sequence, denoted as $\textsc{BI}_{\tau}$, which maximizes ^$\hat{\mathcal{H}}(O_{\tau})$; in other words,
@@
The $\textsc{B}\tau\textsc{I}$ algorithm first construct the ordered sequence of best-$\tau$ items, $\textsc{BI}_{\tau}$, and then extends this sequence to include the remaining items in $U$. The extension can be done using any arbitrary ordering for the remaining items, as it does not affect the algorithm’s approximation ratio. However, a more refined approach, such as a greedy heuristic, could be used to append items that maximize the marginal gain on ^$\hat{\mathcal{H}}(O_{\tau})$ at each step.
The performance of $\textsc{B}\tau\textsc{I}$ algorithm is as follows.
@@
Assume that $p_{i}=p<1$, for all $i\in U$. For any integer $\tau$, where $2\leq\tau\leq n$, the $\textsc{B}\tau\textsc{I}$ algorithm obtains a $(1-p^{\tau-1}-p^{n-\tau}+p^{n})$-approximation for MaxOHP.
Notice that when there is a constant $\epsilon>0$ such that $p<1-\epsilon$, and $\tau\ll n$, both $p^{n-\tau}$ and $p^{n}$ become negligible. In this case, $\textsc{B}\tau\textsc{I}$ is a $\left(1-p^{\tau-1}-\Theta(p^{n})\right)$-approximation algorithm. This aligns with our intuition: optimizing the dominant term $\mathcal{H}(O_{\tau})$ yields a good approximation for the MaxOHP problem.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Information retrievalccs: Theory of computation Approximation algorithms analysis
@@ 1. Introduction
The model needs to be instantiated with a diversity measure for item sets, and any standard measure can be used. In this paper, we consider two commonly-used diversity measures: pair-wise sum diversity (Borodin et al., 2012) and coverage diversity (Ashkan et al., 2015; Puthiya et al., 2016; Zhai et al., 2015). For coverage diversity, we prove the problem is ordered-submodular (Kleinberg et al., 2023), and thus, the standard greedy algorithm achieves a 1/2-approximation.
@@
Depending on the values of the continuation probabilities, we introduce two distinct algorithms. When continuation probabilities are relatively small, e.g., constant values that are smaller than $1$, we design an algorithm, $\small\textsc{B}\tau\textsc{I}$, that focuses on selecting the top-$\tau$ items, where the value of $\tau$ controls the trade-off between runtime and approximation quality. Notably, when $\tau=2$, our approach aligns with a simple greedy approach, and hence, we provide the approximation ratio for this greedy approach. In contrast, when continuation probabilities are arbitrarily close to $1$, it’s not just the first few items that affect the sequential diversity score; rather, more items ranked toward the front impact the score. To address this case, we propose a second algorithm using a greedy matching approach on the set of all items, and offers a constant approximation guarantee.
In this journal extension, we propose two new greedy algorithms that directly tackle the MaxSSD problem. We show that when the continuation probabilities are uniform, i.e., all items share the same probability, ignoring the probabilities and greedily constructing rankings that maximize pairwise distance sums yields a $\frac{1}{2}$-approximation. For non-uniform probabilities, we propose an algorithm, $\small\textsc{SB}\tau\textsc{I}$, that provides a constant-factor approximation guarantee by selecting the best $\tau$ items to form a length-$\tau$ ordering that maximizes the sequential sum over all size-$\tau$ rankings.
Our experiments demonstrate the superiority of our proposed algorithms over the baselines. In particular, the $\small\textsc{SB}\tau\textsc{I}$ algorithms generally achieve better results than the $\small\textsc{B}\tau\textsc{I}$ algorithms, consistent with our theoretical analysis. In addition to the $\mathcal{S}_{+}$ values, we evaluate user satisfaction and engagement metrics for the rankings generated by each approach. Due to space limitations, complete proofs and extensive experimental results are provided in the appendices.
@@ 3.1. Problem Definition
We will work with two commonly-used diversity functions. The first is the pair-wise sum diversity: for a distance function $d\colon U\times U\rightarrow\mathbb{R}_{\geq 0}$, the diversity function $\mathcal{D}_{+}(A)=\sum_{i,j\in A}d\left(i,j\right)$ sums all pair-wise distances of items in $U$.
@@
(3.1) $\mathcal{S}_{+}(O)=\mathbb{E}_{A\sqsubseteq O}[\mathcal{D}_{+}(A)]=\mathbb{E}_{A\sqsubseteq O}\left[\sum\nolimits_{i,j\in A}d\left(i,j\right)\right].$
@@ 3.2. Reformulation and Complexity
To simplify our calculations, we use the notation $p_{O_{i}}=\prod_{t=1}^{i}p_{\pi(t)}$ and $d\left(\pi(i+1),O_{i}\right)=\sum\nolimits_{t=1}^{i}d\left(\pi(i+1),\pi(t)\right)$.
@@
(3.5) $\mathcal{S}_{+}(O)=\sum\nolimits_{i=1}^{n-1}p_{O_{i+1}}d\left(\pi(i+1),O_{i}\right).$
Proof.
We start with demonstrating that the sequential sum-diversity objective can be precisely reformulated via the expression $2\sum_{i=2}^{n}\sum_{j=1}^{i-1}\prod_{t=1}^{i}\mathrm{Pr}_{\pi(t)}d\left(\pi(i),\pi(j)\right)$. Next, the formulation can be simplified using the definition of $p_{O_{i}}$ and $d\left(\pi(i),O_{i-1}\right)$.
(3.6) $\sum_{k=1}^{n}\mathrm{Pr}(A=O_{k})\sum_{\pi(i),\pi(j)\in O_{k}}d\left(\pi(i),\pi(j)\right)\stackrel{{\scriptstyle(a)}}{{=}}\sum_{k=1}^{n}\mathrm{Pr}(A=O_{k})\sum_{i,j\leq k}d\left(\pi(i),\pi(j)\right)$
$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{k=1}^{n-1}(\prod_{t=1}^{k}p_{\pi(t)}-\prod_{t=1}^{k+1}p_{\pi(t)})\sum_{i,j\leq k}d\left(\pi(i),\pi(j)\right)+\prod_{t=1}^{n}p_{\pi(t)}\sum_{i,j\leq n}d\left(\pi(i),\pi(j)\right)$
$=\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(-\sum_{i,j\leq k}d\left(\pi(i),\pi(j)\right)+\sum_{i,j\leq k+1}d\left(\pi(i),\pi(j)\right))$
$=\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(\sum_{i=k+1,j\leq k}d\left(\pi(i),\pi(j)\right)+\sum_{j=k+1,i\leq k}d\left(\pi(i),\pi(j)\right))$
$=2\sum_{k=1}^{n-1}\prod_{t=1}^{k+1}p_{\pi(t)}(\sum_{i=k+1,j\leq k}d\left(\pi(i),\pi(j)\right))=2\sum_{k=2}^{n}\prod_{t=1}^{k}p_{\pi(t)}(\sum_{i=k,j\leq k-1}d\left(\pi(i),\pi(j)\right))$
$=2\sum_{i=2}^{n}\sum_{j=1}^{i-1}\prod_{t=1}^{i}p_{\pi(t)}d\left(\pi(i),\pi(j)\right).$
@@
Let $U=\{u_{1},u_{2},u_{3}\}$, where $d\left(u_{1},u_{2}\right)=0.3$, $d\left(u_{1},u_{3}\right)=1$ and $d\left(u_{2},u_{3}\right)=1$. Let $p_{u_{1}}=p_{u_{2}}=1$ and $p_{u_{3}}=0$. The sequential sum diversity scores differ according to the order of the items, specifically:
@@ 4. Sequential Coverage Diversity
where $||$ denotes the concatenation of sequences and elements.
@@ 5. Ordered Hamiltonian Path
We are given a finite set $U=\{1,\ldots,n\}$ of $n$ distinct items, a distance function $d\left(\cdot,\cdot\right)$, and probabilities $\{p_{1},\ldots,p_{n}\}$ assigned to each item $i\in U$. Let $O=(\pi(i))_{i=1}^{n}$ be the ordered sequence of items of $U$ according to an order $\pi$. The ordered Hamiltonian path, denoted by $\mathcal{H}$, is defined as
(5.1) $\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right),$
@@
Note that the ordered Hamiltonian-path differs significantly from the classic Hamiltonian-path (Gurevich and Shelah, 1987), where the edge weights do not depend on the order in which edges are placed. In contrast, in the ordered Hamiltonian-path, the edge weights $d\left(\pi(i),\pi(i+1)\right)$ are multiplied by a coefficient $W_{O_{i}}$, which depends on the order $\pi$ as well as on the probabilities of all the items in $U$.
@@
We are given a finite set $U=\{1,\ldots,n\}$ of $n$ distinct items, a distance function $d\left(\cdot,\cdot\right)$, and probabilities $p_{1},\ldots,p_{n}$ assigned to each item $i\in U$. The goal is to find an order $O^{*}$ of the items in $U$ so as to maximize the $\mathcal{H}$ objective, that is,
@@
Lemma 0.
$\mathcal{H}(O)$ can be equivalently formulated as
$\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}p_{O_{i+1}}d_{L}(O_{i+1}),$
where $d_{L}(O_{k})\coloneqq\sum_{t=1}^{k-1}d\left(\pi(t),\pi(t+1)\right)$.
We observe that $\mathcal{H}(O)$ can be reformulated according to Lemma 2. The reformulation helps us to relate $\mathcal{H}(O)$ with $\mathcal{S}_{+}(O)$. Notice that $p_{O_{i+1}}$ is the common coefficient of $d_{L}(O_{i+1})$ in $\mathcal{H}(O)$ and $d\left(\pi(i+1),O_{i}\right)$ in $\mathcal{S}_{+}(O)$. Hence, we essentially only compare $d_{L}(O_{i+1})$ with $d\left(\pi(i+1),O_{i}\right)$, for all $1\leq i\leq n-1$.
The following corollary is a direct implication from the triangle inequality, $2d\left(\pi(i+1),O_{i}\right)\geq d_{L}(O_{i+1})$, for any sequence $O$.
Corollary 0.
For any ordered sequence $O$ of items in $U$ it holds
$2\mathcal{S}_{+}{(O)}\geq\mathcal{H}(O).$
Let $O^{*}$ be the optimal sequence of $\mathcal{S}_{+}$: $O^{*}=\arg\allowbreak\max_{O=\pi(U)}\mathcal{S}_{+}{(O)}$. Let $(\pi(i),\pi(i+1))$ be any pair of items swapped in $O^{*}$, and denote $O^{\prime}=(\ldots,\pi(i+1),\pi(i),\ldots)$. By optimality, it follows that $\mathcal{S}_{+}{(O^{*})}\geq\mathcal{S}_{+}{(O^{\prime})}$. With some simplification, we get the following lemma.
Lemma 0.
$d\left(\pi(i+1),O_{i}^{*}\right)\leq\frac{1-p_{\pi(i+1)}}{p_{\pi(i+1)}}\sum\nolimits_{j=1}^{i}\frac{p_{\pi(j+1)}}{1-p_{\pi(j+1)}}d\left(\pi(j),\pi(j+1)\right),$
for any $1\leq i\leq n-1$.
We observe that $\frac{1-p_{i}}{p_{i}}\leq\frac{1-a}{a}$, and $\frac{p_{i}}{1-p_{i}}\leq\frac{b}{1-b}$ if $p_{i}\in[a,b]$. The following corollary holds directly by the implication of Lemma 4 that $d\left(\pi(i+1),O_{i}^{*}\right)\leq\frac{b(1-a)}{a(1-b)}d_{L}(O_{i+1}^{*})$.
Corollary 0.
@@
Let $p_{i}\in[a,b]$, with $0<a<b<1$. An $\alpha$-approximation solution for MaxOHP is a $\frac{a(1-b)}{2b(1-a)}\alpha$-approximation for MaxSSD.
Proof.
where inequality $(\dagger)$ follows from Corollary 5, and $(\ddagger)$ from Corollary 3. Thus we proved $\frac{\mathcal{S}_{+}{(O^{\alpha})}}{\mathcal{S}_{+}{(O^{*})}}\geq\frac{a(1-b)}{2b(1-a)}\alpha$. We note that for uniform continuation probabilities, i.e., $a=b$, $\frac{2b(1-a)}{\alpha a(1-b)}=1/2$. ∎
@@ 6. Uniform Continuation Probabilities
We analyze three regimes for $p$, and design algorithms that achieve constant-factor approximations for MaxSSD in each regime. Due to space constraints, two regimes are presented in the main content, with the third deferred to the full version of this paper.
As outlined earlier, we reformulate the problem using the intemediate problem MaxOHP, which then leads to a solution for MaxSSD. Recall that for an order $O$, the ordered Hamiltonian-path objective is defined as $\mathcal{H}(O)=\sum_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right)$, where $W_{O_{i}}=\sum_{j=i+1}^{n}p_{O_{j}}$. In the uniform probability setting, since $p_{i}=p$ for all $i$, the coefficient $W_{O_{i}}$ is entirely determined by its position in the order, rather than the actual items being ranked. In this case,
@@ 6.1. The best-$\tau$ items Algorithm
To obtain a better intuition for this regime, consider a scenario where $p$ is very small, i.e., close to 0. In this case, a user is likely to quit after examining only a few items, so the ranking of the first few items becomes critical. To maximize diversity, the most diverse items should be placed at the top, as they contribute most significantly to the objective of the ordered Hamiltonian path.
Building on this insight, we propose an algorithm for the general case where $p$ is bounded away from 1. We call this algorithm best-$\tau$ items ($\small\textsc{B}\tau\textsc{I}$), where $\tau$ is a parameter controlling the approximation quality. The algorithm always outputs a ranking of all items, with the pseudocode provided in Algorithm 1.
@@
Algorithm 1 Best-$\tau$ items ($\small\textsc{B}\tau\textsc{I}$) algorithm
Concretely, the best-$\tau$ items ($\small\textsc{B}\tau\textsc{I}$) algorithm works as follows: for any integer $\tau$, let
$\hat{\mathcal{H}}(O_{\tau})=\sum\nolimits_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)$
denote the contribution of $O_{\tau}$ to $\mathcal{H}(O)$. The $\small\textsc{B}\tau\textsc{I}$ algorithm seeks the optimal $\tau$-item sequence, denoted as $\textsc{BI}_{\tau}$, which maximizes $\hat{\mathcal{H}}(O_{\tau})$; in other words,
@@
The $\small\textsc{B}\tau\textsc{I}$ algorithm first construct the ordered sequence of best-$\tau$ items, $\textsc{BI}_{\tau}$, and then extends this sequence to include the remaining items in $U$. The extension can be done using any arbitrary ordering for the remaining items, as it does not affect the algorithm’s approximation ratio. However, a more refined approach, such as a greedy heuristic, could be used to append items that maximize the marginal gain on $\hat{\mathcal{H}}(O_{\tau})$ at each step.
The performance of $\small\textsc{B}\tau\textsc{I}$ algorithm is as follows.
@@
Assume that $p_{i}=p<1$, for all $i\in U$. For any integer $\tau$, where $2\leq\tau\leq n$, the $\small\textsc{B}\tau\textsc{I}$ algorithm obtains a $(1-p^{\tau-1}-p^{n-\tau}+p^{n})$-approximation for MaxOHP.
Proof.
$\mathcal{H}(O)=\sum\nolimits_{i=1}^{n-1}W_{O_{i}}d\left(\pi(i),\pi(i+1)\right).$
$\mathcal{H}(O^{o})=\sum\nolimits_{i=1}^{n-1}W_{O_{i}^{o}}d\left(\pi^{o}(i),\pi^{o}(i+1)\right).$
$\sum_{j=i+1}^{n}p^{j}$ $\stackrel{{\scriptstyle(a)}}{{=}}p^{i+1}\frac{1-p^{n-i}}{1-p}=\frac{p^{i+1}}{1-p}-\frac{p^{n+1}}{1-p},$
where equality $(a)$ holds by Lemma 1.
Recall that the $\small\textsc{B}\tau\textsc{I}$ algorithm finds the optimal $\tau$-item sequence that maximizes the following equation:
(6.3) $O_{\tau}=\operatorname*{arg\,max}_{O_{\tau}\sqsubseteq\pi(U)}\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right).$
For the convenience in our calculations, we let
$\ell_{\pi}=\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right).$
Let us derive a lower bound on $\mathcal{H}(O)$ as follows:
(6.4) $\mathcal{H}(O)$ $=\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)-\sum_{i=1}^{n-1}\frac{p^{n+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)$
$\geq\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)-\sum_{i=1}^{\tau-1}\frac{p^{n+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)$
$=\ell_{\pi}-\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1)).$
(6.5) $\mathcal{H}(O^{o})$ $=\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)-\sum_{i=1}^{n-1}\frac{p^{n+1}}{1-p}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)$
$\leq\sum_{i=1}^{n-1}\frac{p^{i+1}}{1-p}d\left(\pi^{o}(i),\pi^{o}(i+1)\right)\leq\sum_{t=1}^{T}\sum_{i=(t-1)(\tau-1)+1}^{t\cdot(\tau-1)}\frac{p^{i+1}}{1-p}d(\pi^{o}(i),\pi^{o}(i+1))$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\ell_{\pi}(1+p^{\tau-1}+\cdots+p^{(T-1)(\tau-1)})\leq\frac{\ell_{\pi}}{1-p^{\tau-1}}.$
Combining Equations (6.4) and (6.5), we get
$=(1-p^{\tau-1})\left(1-\frac{\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1))}{\sum_{i=1}^{\tau-1}\frac{p^{i+1}}{1-p}d\left(\pi(i),\pi(i+1)\right)}\right)$
$\geq(1-p^{\tau-1})\left(1-\frac{\frac{p^{n+1}}{1-p}\sum_{i=1}^{\tau-1}d(\pi(i),\pi(i+1))}{\sum_{i=1}^{\tau-1}\frac{p^{\tau}}{1-p}d\left(\pi(i),\pi(i+1)\right)}\right)$
$=(1-p^{\tau-1})(1-p^{n-\tau+1})=1-p^{\tau-1}-p^{n-\tau+1}+p^{n}\geq 1-p^{\tau-1}-p^{n-\tau}+p^{n}.$
We have demonstrated that the $\small\textsc{B}\tau\textsc{I}$ algorithm provides an asymptotic $(1-p^{\tau-1}-p^{n-\tau}+p^{n})$-approximation. ∎
Notice that when there is a constant $\epsilon>0$ such that $p<1-\epsilon$, and $\tau\ll n$, both $p^{n-\tau}$ and $p^{n}$ become negligible. In this case, $\small\textsc{B}\tau\textsc{I}$ is a $\left(1-p^{\tau-1}-\Theta(p^{n})\right)$-approximation algorithm. This aligns with our intuition: optimizing the dominant term $\mathcal{H}(O_{\tau})$ yields a good approximation for the MaxOHP problem.